OpenStudy (anonymous):

The cost of a ticket to an amusement park was increased by 4 dollars. The new cost is twenty-eight dollars. a. Write a one-step equation that could be solved to find the old price. Let p represent the old price. b. Solve the equation you wrote in Part a. Show all your work. c. Explain how you know that the solution you found in Part b is the correct answer.

OpenStudy (mathstudent55):

If p is the old price, and the new price is $4 more, how would you write that?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Let me ask you this way. Let p represent a number. How do you write 4 more than p, or 4 added to p?

OpenStudy (anonymous):

by adding to both sides?

OpenStudy (mathstudent55):

There are no sides here. We don't have an equation yet. An equation will come soon. Think of the number 20. Without doing the addition, how do you write 4 more than 20, or 4 added to 20? Don't add 20 and 4, just show the addition.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

In this last question, what I was looking for was 20 + 4 That shows the addition of 20 and 4, or 4 more than 20. In your problem above, it's similar. You have a number represented by the variable p. You need to represent 4 more than p. You do that by writing p + 4 The old price is p. The new price is 4 more than p, so it's p + 4. Now we get to the equation. We are told the new price is $28. Since we know the new price is p + 4, and the new price is also $28, that means p+4 must be the same a 28. Now we write an equation: p + 4 = 28

OpenStudy (anonymous):

ight

OpenStudy (anonymous):

now i add 4 to both sides

OpenStudy (anonymous):

?

OpenStudy (mathstudent55):

This is the answer to part a. p + 4 = 28

OpenStudy (mathstudent55):

In part b. you need to solve the equation.

OpenStudy (anonymous):

p is 24

OpenStudy (anonymous):

@m

OpenStudy (mathstudent55):

Correct. The way to solve the equation is that since you have p + 4 on the left side and you want p alone, you must get rid of the plus 4. The way you get rid of an addition is by doing the opposite operation which is subtraction. p + 4 - 4 = p In an equation, you must do the same operation to both sides, so you must subtract 4 from both sides. p + 4 - 4 = 28 - 4 p = 24 That is the solution to part b.

OpenStudy (anonymous):

ok so wats c.

OpenStudy (mathstudent55):

Now for part c. you must explain why you know the solution p = 24 is correct. Since the problem stated the old price plus 4 is 28, we can try it with 24 to see if it works. Is 24 + 4 equal to 28? The answer is yes, 24 + 4 equals 28, so our solution is correct. Not only that, the only number there is that when 4 is added to it will give you 28 is 24. There is no other number that can be the solution.

OpenStudy (anonymous):

ight thx 4 ur help

OpenStudy (mathstudent55):

you're wlcm